OpenStudy (anonymous):

Integral of x^(1/2)*e^(x^(1/2)), using Integration by parts and substitution, if needed. It doesn't seem like substituting x^1/2 would work, cause there isn't anything close to (x^(-1/2)/2), but I'm not doing great just integrating it out...

OpenStudy (freckles):

\[\int\limits \frac{x e^{x^\frac{1}{2}}}{x^\frac{1}{2}} dx =2 \int\limits \frac{x e^{x^\frac{1}{2}}}{2x^\frac{1}{2}} dx \\ \text{ Let } u=x^\frac{1}{2}\]

OpenStudy (freckles):

once your write the integral in terms of u the integration by parts should be easy to see